21:02 last night

Date view Thread view Subject view Author view Attachment view

From: W.J. Llope (llope_at_physics.rice.edu)
Date: Wed Jan 22 2003 - 09:28:31 EST


from the shift log Jan 21, 2003 21:02:

21:02 2 issues - No beam yet this shift, and no estimate. The MCR monitor
states 'machine setup', instead of 'tandem fault'.

Bill Christie called earlier. He requested a normal run (TOF in) and
a short test run with TOF out.
Histograms should be saved hardcopy for these two runs. There is
concern about problems
similar to last year involving the TOF, and these data are to test if
the problems are solved.

- Hal Spinka

dear bill ch,
we request that you contact tofp_at_physics.rice.edu
if you have any questions/concerns/comments about tofp operations...
a cryptic note in the shift log is not sufficient communication
if there is a problem.... please see
http://mac8.rice.edu/~TOF/TOFp/operations/tofp_shiftcrewFYI.html

now, what problem are you talking about? the problem last run
when TRG changed the L0 timing whout telling anyone?!? LOL

this run, we've been talking good data (2.5M events by my count) -
is there a new issue?

i will be trying to call you throughout the morning

-- 

W.J. Llope, Ph.D. Res. Assoc. Professor T.W. Bonner Nuclear Laboratory phone: 713-348-4741 Rice University, MS-315 fax: 713-348-5215 6100 S. Main St. Houston, TX 77005-1892 e-mail: llope_at_physics.rice.edu WWW home: http://wjllope.rice.edu/default.html

--


Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:32 EDT